Hi
A had 52 dollars more than B.B had 3 times as much as C. If C had 100 dollars less than A how much money did they have altogether.
Thank you

Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

let a equal the amount that A had.
let b equal the amount that B had.
let C equal the amount that C had.

yu are given that:

a = b + 52
b = 3c
c = a - 100

in the first 2 equations, replace b with 3c to get:
a = b + 52 becomes a = 3c + 52

in the third equation, replace a with 3c + 52 to get:
c = a - 100 becomes c = 3c + 52 - 100.
combine like terms to get:
c = 3c - 48
subtract c from both sides of the equation and add 48 to both sides of the equation to get:
48 = 2c
solve for c to get:
c = 24

in the equation that says b = 3c, replace c with 24 to get:
b = 3 * 24.
solve for b to get:
b = 72.

in the equation that says a = b + 52, replace b with 72 to get:
a = 72 + 52 = 124.

your now have:

a = 124
b = 72
c = 24

add then up and you have a + b + c = 220.

that's your solution.



Answer by ikleyn(52832)

.

Let C = x;

then B = 3x  and  A = 3x+52.


You are given  


    A - C = 100,  or

    (3x + 52) - x = 100,

     2x           = 100-52 = 48

      x                    = 48/2 = 24.


C had 24 dollars;  B had  3*24 = 72 dollars;  A had 72 + 52 = 124 dollars.


Together, they had   24 + 72 + 124 = 220 dollars.      ANSWER

Solved.

=================

There is special  SHARM,  when the problem solution is presented in a short form.

As short,  as it  SHOULD  be.


This sharm is called  MATHEMATICAL  STYLE.